OpenStudy (anonymous):

Find the difference quotient and simplify your answer. f(x)=((4)/(x+1), f(x)-f(7))/x-7, x≠7

OpenStudy (loser66):

f(7)=?

OpenStudy (anonymous):

i goingtogive you more clear

OpenStudy (anonymous):

([(4)/(x+1)]-[(4)/7+1)])/(x-7)

OpenStudy (loser66):

\[\dfrac{\frac{4}{x+1}-\frac{4}{8}}{x-7}\]right?

OpenStudy (anonymous):

yeah like that

OpenStudy (loser66):

solve the big numerator first. what do you get? (Hint: make them have the same denominator). I mean